# OneOcc: Semantic Occupancy Prediction for Legged Robots with a Single Panoramic Camera

### v1

@@ 3.5 Gait Displacement Compensation
where $\mathrm{Linear}_{0}$ has all weights and bias set to zero so the initial warp is identity (cf. zero-conv [zhang2023adding]). Before lifting we correct and resample:
@@
Here, $v\!\in\!\{\text{unwrapped},\text{spherical}\}$ indexes the projection path; $\pi_{v}(\mathbf{c};\boldsymbol{\kappa})$ projects a 3D voxel centroid $\mathbf{c}$ to image pixels using the calibrated Taylor camera with parameters $\boldsymbol{\kappa}$. $\widehat{\mathbf{p}}$ is the corrected sampling pixel; $\mathrm{bilinear}(\cdot)$ samples the 2D feature map $\mathbf{F}^{v}_{1/s}$, and $\mathbf{V}^{(\cdot)}_{s}$ denotes the lifted 3D feature.
@@
GDC routes the phase error back to 2D, preventing mis-attributed voxels. The zero-init head stabilizes early training and leaves features unperturbed when motion is tiny ($\Delta_{s}\!\approx\!0$).
@@ 3.6 Hierarchical AMoE-3D
Method.
The 3D decoder is a depthwise-separable UNet with trilinear upsampling at three levels (L1/L2/L3). At each level we fuse $\widetilde{\mathbf{V}}^{\mathrm{Ca}}$ using a dual-path volumetric saliency and a Mixture-of-Experts (MoE) fuse; we term this module AMoE-3D, where A denotes attention implemented via the dual-path (channel & spatial) saliency gates.
@@ 3.7 Segmentation Head and Loss
We follow MonoScene [cao2022monoscene] for most losses: (i) standard cross-entropy on valid voxels, (ii) the scene-class affinity (SCAL) terms for semantics and geometry, and (iii) the frustums proportion (FP) loss. We do not use the relation loss, as we observed it encourages co-occurrence priors that over-smooth azimuthal boundaries and suppress small near-field classes under panoramic legged setting, yielding no gain and occasional degradation on our data. The final objective is:
@@ 4 Experiments
SSCNet [song2017semantic] 67.12 62.95 48.12 14.03 52.97 76.34 45.50 14.88 57.81 75.30 48.60 9.51
SSCNet-full [song2017semantic] 80.56 58.52 51.28 16.04 73.79 64.34 52.37 17.14 73.64 67.31 54.24 11.56
LMSCNet [roldao2020lmscnet] 81.18 62.50 54.59 17.33 75.34 70.92 57.56 18.91 75.68 72.67 58.91 13.40
OccRWKV [wang2024occrwkv] 72.84 63.83 51.56 16.25 66.43 59.30 45.63 15.42 74.10 58.31 48.44 12.61
@@ 4.1 Setup
Datasets. We evaluate on two panoramic occupancy benchmarks for legged/humanoid platforms. QuadOcc: real first-person $360^{\circ}$ data on a quadruped within a campus; day/dusk/night, sequence-heterogeneous split; 2424K consecutive frames with stride-5 training; $6$ classes on a $64{\times}64{\times}8$ grid (0.4 m). Human360Occ (H3O): CARLA-based human-ego $360^{\circ}$ with simulated gait over $16$ maps and diverse weather/lighting; RGB, depth, occupancy (two resolutions), and poses. We report within-city (H3O-Homo) and cross-city (H3O-Heter). See Table 2 and Table 3.
@@
Baselines. We re-train popular LiDAR SSC systems (SSCNet [song2017semantic], LMSCNet [roldao2020lmscnet], OccRWKV [wang2024occrwkv]) and adapt representative vision SSC methods (OccFormer [zhang2023occformer], VoxFormer [li2023voxformer], SGN [mei2023camera], MonoScene [cao2022monoscene]) to panoramas via calibrated unwrapping to equirectangular. For P+L entries in Table 2, we add LiDAR to the corresponding vision model per the table footnote.
Implementation. Unless noted, training follows official configs with minimal dataset-specific changes; schedules/image sizes/voxelization match for fairness. On H3O we disable the raw-projection branch (native equirectangular), keeping components as in ablations. Training uses a single NVIDIA RTX-4090 GPU.
@@ 4.3 Ablations
We ablate GDC, DP-ER, BGV, AMoE-3D (Table 4). Starting from Q0 (ER-only, no GDC, single-grid, no AMoE-3D; 19.19 mIoU), we add: +GDC $\rightarrow$ 19.58 (+0.39), +DP-ER $\rightarrow$ 19.89 (+0.31), +BGV $\rightarrow$ 20.30 (+0.41), +AMoE-3D (full) $\rightarrow$ 20.56 (+0.26). Gains are additive: GDC stabilizes sampling; DP-ER provides complementary cues; BGV reduces discretization bias; AMoE-3D sharpens edges/contacts without over-smoothing flats.
@@ 4.5 Efficiency
Inference on an RTX 4090 at $608{\times}1216$, batch size $1$. OneOcc averages 69.93 ms (FPS $\approx$ 14.30). The model has 101.76M parameters and peaks at 1.86 GB CUDA allocation (2.65 GB reserved) in FP32. With mixed precision (FP32+FP16), latency drops to 52.84 ms (FPS $\approx$ 18.92), and peak allocation is 1.49 GB. The footprint remains moderate, making OneOcc suitable for onboard legged perception on a single GPU. Jetson Thor extrapolation: using published tensor-core budgets (RTX 4090 FP16 tensor 165 TFLOPS vs. Thor 517 TFLOPS at FP16 with 2:4 sparsity), a linear scale suggests >>10 FPS is achievable on Thor.
@@ 5 Conclusion
Limitations & Outlook. OneOcc assumes accurate calibration with bounded drift. A possible remedy is online extrinsic self-calibration, optionally regularized by lightweight odometry priors. We propose robotic occupancy as a intermediate layer for world models and vision-language-action models: tokens for language control, and occupancy-sequence pretraining for cross-robot transfer.

### v2

@@ 3.5 Gait Displacement Compensation
where $\mathrm{Linear}_{0}$ has all weights and bias set to zero so the initial warp is identity (cf. zero-conv [105]). Before lifting we correct and resample:
@@
In practice, this also upgrades the lifting operator from integer-index gather [4] to bilinear sampling, which reduces projection aliasing and further stabilizes GDC. Here, $v\!\in\!\{\text{unwrapped},\text{spherical}\}$ indexes the projection path; $\pi_{v}(\mathbf{c};\boldsymbol{\kappa})$ projects a 3D voxel centroid $\mathbf{c}$ to image pixels using the calibrated Taylor camera with parameters $\boldsymbol{\kappa}$. $\widehat{\mathbf{p}}$ is the corrected sampling pixel; $\mathrm{bilinear}(\cdot)$ samples the 2D feature map $\mathbf{F}^{v}_{1/s}$, and $\mathbf{V}^{(\cdot)}_{s}$ denotes the lifted 3D feature.
@@
GDC routes the phase error back to 2D, preventing mis-attributed voxels. The zero-init head stabilizes early training and leaves features unperturbed when motion is tiny ($\Delta_{s}\!\approx\!0$). We additionally provide qualitative analysis of GDC in the supplementary material.
@@ 3.6 Hierarchical AMoE-3D
Method. The 3D decoder is a depthwise-separable UNet with trilinear upsampling at three levels (L1/L2/L3). At each level we fuse $\widetilde{\mathbf{V}}^{\mathrm{Ca}}$ using a dual-path volumetric saliency and a Mixture-of-Experts (MoE) fuse; we term this module AMoE-3D, where A denotes attention implemented via the dual-path (channel & spatial) saliency gates.
@@ 3.7 Segmentation Head and Loss
We follow MonoScene [4] for most losses: (i) standard cross-entropy on valid voxels, (ii) the scene-class affinity (SCAL) terms for semantics and geometry, and (iii) the frustums proportion (FP) loss. We do not use the relation loss, as we observed it encourages co-occurrence priors that over-smooth azimuthal boundaries and suppress small near-field classes under panoramic legged setting, yielding no gain and occasional degradation on our data. The final objective is:
@@ 4 Experiments
SSCNet [76] 67.12 62.95 48.12 14.03 52.97 76.34 45.50 14.88 57.81 75.30 48.60 9.51
SSCNet-full [76] 80.56 58.52 51.28 16.04 73.79 64.34 52.37 17.14 73.64 67.31 54.24 11.56
LMSCNet [71] 81.18 62.50 54.59 17.33 75.34 70.92 57.56 18.91 75.68 72.67 58.91 13.40
OccRWKV [85] 72.84 63.83 51.56 16.25 66.43 59.30 45.63 15.42 74.10 58.31 48.44 12.61
@@ 4.1 Setup
Datasets. We evaluate on two panoramic occupancy benchmarks for legged/humanoid platforms. QuadOcc: real first-person $360^{\circ}$ data on a quadruped within a campus; day/dusk/night, sequence-heterogeneous split; 2424K consecutive frames with stride-5 training; $6$ classes on a $64{\times}64{\times}8$ grid (0.4 m). Human360Occ (H3O): CARLA-based human-ego $360^{\circ}$ with simulated gait over $16$ maps and diverse weather/lighting; RGB, depth, occupancy (two resolutions), and poses. We report within-city (H3O-Homo) and cross-city (H3O-Heter). See supplementary for the data collection pipelines and statistics of QuadOcc and H3O.
@@
Baselines. We re-train popular LiDAR SSC systems (SSCNet [76], LMSCNet [71], OccRWKV [85]) and adapt representative vision SSC methods (OccFormer [107], VoxFormer [43], SGN [56], MonoScene [4]) to panoramas via calibrated unwrapping to equirectangular. For P+L entries in Table 2, we add LiDAR to the corresponding vision model per the table footnote. Details in the supplementary.
Implementation. Unless noted, training follows official configs with minimal dataset-specific changes; schedules/image sizes/voxelization match for fairness. On H3O we disable the raw-projection branch (native equirectangular), keeping components as in ablations. Training uses a single RTX-4090 GPU. Details in the supplementary.
@@ 4.3 Ablations
We ablate GDC, DP-ER, BGV, AMoE-3D (Table 4). Starting from Q0 (ER-only, no GDC, single-grid, no AMoE-3D; 19.19 mIoU), we add: +GDC $\rightarrow$ 19.58 (+0.39), +DP-ER $\rightarrow$ 19.89 (+0.31), +BGV $\rightarrow$ 20.30 (+0.41), +AMoE-3D (full) $\rightarrow$ 20.56 (+0.26). Gains are additive: GDC stabilizes sampling; DP-ER provides complementary cues; BGV reduces discretization bias; AMoE-3D sharpens edges/contacts without over-smoothing flats. We further provide more detailed per-module ablations, calibration-robustness analysis, temporal aggregation baselines, and resolution scaling results in the supplementary material.
@@ 4.5 Efficiency
Inference on an RTX 4090 at $608{\times}1216$, batch size $1$. OneOcc averages 69.93 ms (FPS $\approx$ 14.30). The model has 101.76M parameters and peaks at 1.82 GB CUDA allocation (2.58 GB reserved) in FP32. With mixed precision (FP32+FP16), latency drops to 52.84 ms (FPS $\approx$ 18.92), and peak allocation is 1.49 GB. The footprint remains moderate, making OneOcc suitable for onboard legged perception on a single GPU. We additionally report measured on-device runtime and a module-wise latency breakdown on NVIDIA Jetson AGX Orin in the supplementary material.
@@ 5 Conclusion
Limitations & Outlook. OneOcc assumes accurate calibration with bounded drift. A possible remedy is online extrinsic self-calibration, optionally regularized by lightweight odometry priors. We propose robotic occupancy as an intermediate layer for world models and vision-language-action models: tokens for language control, and occupancy-sequence pretraining for cross-robot transfer.
Acknowledgment
This research was funded by the Natural Science Foundation of Zhejiang Province (Grant No. LZ24F050003), the National Natural Science Foundation of China (Grant Nos. 12174341 and 62473139), the Hunan Provincial Research and Development Project (Grant No. 2025QK3019), the MOE AcRF Tier 1 SSHR-TG Incubator Grant FY24 (Grant No. RSTG7/24), and the opening project of the State Key Laboratory of Autonomous Intelligent Unmanned Systems (Grant No. ZZKF2025-2-10). Hao Shi was also supported by the China Scholarship Council (CSC) for joint Ph.D. training at Nanyang Technological University (NTU).